OpenStudy (anonymous):

let f(x)=x^2+x-3 and g(x)=2x+4. Find g(-1) and f(g(-1))

OpenStudy (jamesj):

Well, g(-1) = -2 + 4 = 2 Agreed? So now f(g(-1)) = f(2) = ....

OpenStudy (anonymous):

yes

OpenStudy (phi):

f(x) is just short hand for x*x + x -3 (because math people don't like to type) every where you see an x, replace it with 2 to get f(2)

OpenStudy (anonymous):

thanks

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

g(-1)=2 and f(g(-1))=3

OpenStudy (anonymous):

is that right?

OpenStudy (phi):

yes

OpenStudy (anonymous):

:) thanks hun